a computer company is purchased $1700. each year , its value is 75% of its value the year before. after how many years will the laptop computer be worth $100 or less?
Answer by fion23(5)

1700(1-r)^1=1700*0.75
1700(1-r)=1275
(1-r)=1275/1700
-r=0.75-1
r=0.25
1700(1-0.25)^t=100
0.75^t=100/1700
t log0.75 = log(1/17)
t=(log(1/17))/(log0.75)
t=9.848 years